OpenStudy (anonymous):

On a subway route, station C is located at the midpoint between stations A and D. Station B is located at the midpoint between stations A and C. If the distance between stations A and D is 2.4 kilometers, what is the distance between stations B and D?

hero (hero):

So first, you have to find the midpoint of A and D, which is C. Then you have to find the midpoint between A and C. A =0 D = 2.4 B = ? C=? C = 2.4/2 = 1.2km B = 1.2/2 = 0.6km

hero (hero):

2.4-0.6 = 1.8km So the distance between B and D is 1.8 km

OpenStudy (agreene):

I would argue this question... but I tend to be that way...

hero (hero):